# Shadow of Quantum Improved Regular Kerr Black Hole and parameter constrains with EHT observations

### v1

@@ III Null geodesics and Photon sphere
$\frac{\partial S}{\partial\tau}=-\frac{1}{2}g^{\alpha\beta}\frac{\partial S}{\partial x^{\alpha}}\frac{\partial S}{\partial x^{\beta}},$ (15)
where $\tau$ is the affine parameter along the geodesics, and $S$ is the Jacobi action. The metric (3) exhibits both time translational and rotational invariance, which leads to conserved quantities along the geodesics: the energy $E=-p_{t}$ and the axial angular momentum $L=p_{\varphi}$, where $p_{\mu}$ represents the photon’s four-momentum. The Petrov-type D nature of the metric (3) guarantees the existence of Carter’s separable constant $\mathcal{K}$, allowing the action to be written in the form:
@@
The “+” and “-” signs in Eq. (19) correspond to outgoing and ingoing photons in the radial direction, while in Eq. (20) they represent photons moving towards the south pole $\theta=\pi$ and the north pole $\theta=0$, respectively. The geodesic equations described above govern the propagation of photons in the QIRK spacetime. When studying photon orbits, two distinct impact parameters are typically defined Chandrasekhar_1985 :
@@
These two impact parameters play a crucial role in determining the shape of a photon’s orbit. Based on an analysis of the radial effective potential $\mathcal{V}_{r}(r)$, photon orbits can be classified into three distinct categories: scattered orbits, absorbed orbits, and bounded orbits (also known as spherical orbits or photon sphere). If these spherical orbits are unstable, small perturbations will cause them to be observable, forming an apparent boundary that separates the dark and bright regions in the observer’s sky, ultimately shaping the optical appearance of the black hole. These orbits are determined by the following conditions:
$R(r)=0\,,\quad\quad\frac{\mathrm{d}R(r)}{\mathrm{d}r}=0\,.$ (24)
@@
If we are considering the rotating black hole, i.e., $a\neq 0$, the solutions of (25) and (26) are
@@
They are all functions of $r_{\text{sphere}}$, i.e. the radius of the photon sphere. At the same time,
@@
If $R^{\prime\prime}(r_{\text{sphere}})>0$, the photon sphere is unstable, while for $R^{\prime\prime}(r_{\text{sphere}})<0$, it is stable.
In the case of a non-rotating black hole ($a=0$), the system exhibits spherical symmetry, allowing us to set $\theta=\pi/2$. In this case, the motion in the $\theta$-direction can be simplified as $\mathcal{K}=0$ (or $\eta=0$). Meanwhile, the impact parameter $\xi$ corresponds to $b$, which is the impact parameter in the spherical symmetric case. Consequently, the equations $R(r)=0$ and $R^{\prime}(r)=0$ reduce to:

### v2

@@ III Null geodesics
$\frac{\partial S}{\partial\lambda}=-\frac{1}{2}g^{\alpha\beta}\frac{\partial S}{\partial x^{\alpha}}\frac{\partial S}{\partial x^{\beta}},$ (15)
where $\tau$ is the affine parameter along the geodesics, and $S$ is the Jacobi master function. The metric (3) exhibits both time translational and rotational invariance, which leads to conserved quantities along the geodesics: the energy $E=-p_{t}$ and the axial angular momentum $L=p_{\varphi}$, where $p_{\mu}$ represents the photon’s four-momentum. The Petrov-type D nature of the metric (3) guarantees the existence of Carter’s separable constant $\mathcal{K}$, allowing the action to be written in the form:
@@
The ”+” and ”-” signs in Eq. (19) correspond to outgoing and ingoing photons in the radial direction, while in Eq. (20) they represent photons moving towards the south pole $\theta=\pi$ and the north pole $\theta=0$, respectively. The geodesic equations described above govern the propagation of photons in the QIRK spacetime. When studying photon orbits, two distinct impact parameters are typically defined Chandrasekhar_1985 :
@@
These two impact parameters play a crucial role in determining the shape of a photon’s orbit. Based on an analysis of the radial effective potential $\mathcal{V}_{r}(r)$, photon orbits can be classified into three distinct categories: scattered orbits, absorbed orbits, and bounded orbits (also known as spherical orbits or photon sphere). If a spherical orbit is unstable, small perturbations allow photons to escape and create the observable photon ring image. These orbits are determined by the following conditions:
$R(r)=0\,,\quad\quad\frac{\mathrm{d}R(r)}{\mathrm{d}r}=0\,,$ (24)
@@
When we consider a nonzero rotation parameter, the solutions of (25) and (26) include an additional family compared to the Kerr case. One such family is
@@
The first family coincides with the Kerr solution; the second is specified by $r=0$ and $\eta=0$. At the same time, it is not hard to find
@@
The photon sphere is unstable when $R^{\prime\prime}(r_{\text{sphere}})>0$ and stable when $R^{\prime\prime}(r_{\text{sphere}})<0$. For the family with $r=0,\ \eta=0$, stability depends on $\xi$: it is unstable when $\lvert\xi\rvert<a$ and stable when $\lvert\xi\rvert>a$.
In the case of a non-rotating black hole ($a=0$), the system exhibits spherical symmetry, allowing us to set $\theta=\pi/2$. In this case, the motion in the $\theta$-direction reduces to $\mathcal{K}=0$ (equivalently $\eta=0$). Meanwhile, the impact parameter $\xi$ corresponds to $b$, which is the impact parameter in the spherical symmetric case. Consequently, the equations $R(r)=0$ and $R^{\prime}(r)=0$ reduce to: